How many grams of CaO will be produced from the decomposition of 1.44 x 10^23 molecules of CaCO3
1
Expert's answer
2020-12-15T07:54:20-0500

CaO is produced by reaction:

CaCO3 = CaO + CO2


One mole of any substance contains 6.022×1023 atoms/molecules (Na).

Hence,

n(CaCO3) = (1.44×1023 molecules of CaCO3) × (1 mol / 6.022×1023 molecules) = 0.239 moles CaCO3


According to the equation: n(CaCO3) = n(CO2)

Thus:

n(CO2) = n(CaCO3) = 0.239 moles


Moles of CaO = Mass of CaO / Molar mass of CaO

The molar mass of CaO is 56.077 g mol-1.

Hence,

Mass of CaO = Moles of CaO × Molar mass of CaO

Mass of CaO = 0.239 mol × 56.077 g mol-1 = 13.4 g

Mass of CaO is 13.4 grams.


Answer: 13.4 grams of CaO